OpenStudy (anonymous):

Algebra 2 help please---Solve the system by using a table. {-0.2x-0.3y=0.2 {0.4x-0.3y=1.4 (2,2) (2,-2) (2,1) (-2,2) {

OpenStudy (anonymous):

start by multiplying both sides of the equation by 10 (for both equations)

OpenStudy (anonymous):

so like i'd multiply -0.2(10)-0.3(10)= -2-3 and 0.4(10)-0.3(10)=4-3? Sorry I'm really bad at this kind of math...

OpenStudy (anonymous):

yeah you're right about the left hand side of both equations. All you gotta do is do the same thing with the right hand side too. Like for the first one, you write 0.2(10)

OpenStudy (anonymous):

I'm so confused...where'd the 10 come from in the first place?

OpenStudy (anonymous):

you're multiplying the whole equation by 10 just to make things easier. It doesn't change the equation in any way. For example, if 3x=4, it doesn't make any difference if you write 3(10)x=4(10) or 30x=40 It's just like that.